Variance Analysis, Revision of Standards, Journal Entries

The Lubbock plant of Morril's Small Motor Division produces a major subassembly for a 6.0 horsepower motor for lawn mowers. The plant uses a standard costing system for production costing and control. The standard cost sheet for the subassembly follows:

Direct materials (6.0 lbs. @ $5) $30.00
Direct labor (1.6 hrs. @ $12) 19.20

During the year, the Lubbock plant had the following actual production activity:

  1. Production of subassemblies totaled 50,000 units.
  2. A total of 260,000 pounds of raw materials was purchased at $4.70 per pound.
  3. There were 60,000 pounds of raw materials in beginning inventory (carried at $5 per lb.). There was no ending inventory.
  4. The company used 82,000 direct labor hours at a total cost of $1,066,000.

The Lubbock plant's practical activity is 60,000 units per year. Standard overhead rates are computed based on practical activity measured in standard direct labor hours.

Required:

1. Conceptual Connection: Compute the materials price and usage variances.

Materials price variance $
Materials usage variance $

Of the two materials variances, which is viewed as the most controllable?

2. Conceptual Connection: Compute the labor rate and efficiency variances.

Labor rate variance $
Labor efficiency variance $

3. Conceptual Connection: Assume that the purchasing agent for the small motors plant purchased a lower-quality raw material from a new supplier. Would you recommend that the plant continue to use this cheaper raw material?

4. Prepare all possible journal entries. For a compound transaction, if an amount box does not require an entry, leave it blank or enter "0".

Journal
Account & Explanation Debit Credit
Record materials price variance
Record materials usage variance
Record payroll
Transfer of Material usage variances and labor variances.
Transfer materials price variance
Answer #1

1) Material Price Variance = (AP – SP)AQ

=(4.70 - 5) 260,000

=78,000 (Favorable)

Material Usage Variance = (AQ – SQ)SP

=(320,000 - 300,000) 5

=100,000 (Unfavorable)

The Market Usage Variance can be controllable more easily if we compare with Material Price variance where the prices are completely dependant on the market which changes and we dont have any control on that. If the proper quality of material is purchased then the usage will less , so for any variance , responsibility will be given to Purchasing.

2) Labor rate variance = (AR – SR)AH

=(13 - 12) 82,000

=82,000 (Unfavorable)

Labor efficiency variance = (AH – SH)SR

=(82,000 - 80,000) 12

=24,000 (Unfavorable)

3)No , we wont recommend the purchasing agent to purchase a lower quality raw material from a new supplier.

Because of the use of Low quality raw material , below is affected variance

Materials price variance $    78,000.00 Favorable
Materials usage variance $ 100,000.00 Unfavorable
Labor efficiency variance $    24,000.00 Unfavorable
Net Effect $    46,000.00 Unfavorable

4)

Materials $ 1,300,000.00
Materials price variance $        78,000.00
Accounts Payable $ 1,222,000.00
Work in Process $ 1,500,000.00
Materials usage variance $      100,000.00
Materials $ 1,600,000.00
Work in Process $      960,000.00
Labor rate variance $        82,000.00
Labor efficiency variance $        24,000.00
Accrued Payroll $ 1,066,000.00
Cost of Goods Sold $      206,000.00
Materials usage variance $      100,000.00
Labor rate variance $        82,000.00
Labor efficiency variance $        24,000.00
Materials price variance $        78,000.00
Cost of Goods Sold $        78,000.00